# Kicking for Goal or Touch? An Expected Points Framework for Penalty Decisions in Rugby Union

### v1

@@ 1 Introduction
While the choice is often straightforward near the opposition’s corner, given a certain point difference, or late in a match, many game states are ambiguous and governed by convention, perceived kicker ability, and the expected quality of a team’s lineout. In contrast, American football has long relied on Expected Points (EP), Expected Points Added (EPA), and win-probability (WP) frameworks to guide fourth-down decision-making, translating field position and context into expected returns and actionable recommendations (Romer2006; Burke2009; YurkoVenturaHorowitz2018). In that setting, WP and EP estimators are typically built using statistical machine learning models (e.g., Random Forests and gradient boosting), as is done in LockNettleton2014 and Baldwin2021.
In rugby union, public research is only beginning to close this methodological gap, which can be largely attributed to a lack of publicly-available data. Martinez-Arastey2025 propose a transparent, phase-level modeling framework for calculating Expected Points on English Premiership data, highlighting the difficulty of reliably predicting minority outcomes (e.g., penalty kicks) and providing a reproducible benchmark rather than a deployable decision tool. Separately, public-facing analyses model goal-kicking success as a function of distance and lateral angle, offering practically-useful (albeit context-free) estimates (Monpezat2023). Building on these strands—and informed by lessons from the NFL literature on decision modeling—we directly evaluate the penalty decision.
@@ 2.1 Phase-Level Data
Our primary dataset consists of phase-level event logs from the 2018/19 Premiership Rugby season compiled by Martinez-Arastey2025, covering 132 matches and 35,199 phases of play. Each record describes the on-field state at the start of a phase: the initiating event type, a zonal field location and side, the current score differential, and the next score in points for or against the team in possession. We harmonize timestamps to “seconds remaining in match” so that time is strictly decreasing within a game and directly comparable across halves.
@@ 2.3 Kicking Data
Our kicking data come from the analysis in QuarrieHopkins2015, which studied 582 international rugby matches played between 2002 and 2011. The resulting dataset consists of 3,802 penalties, aggregated into a 5 m-by-5 m grid over the pitch, with the observed proportion of successful kicks recorded for each grid square. These cell-level success rates serve as the raw input for our continuous angle–distance model of kick success. The structure of this dataset is included in Table 2.
@@ 3.2.1 Dealing with Double Counting
This procedure ensures that each run_id contributes at most one observation to the lineout EP model, mitigating bias from repeated, structurally identical possessions. Our approach is conceptually similar to the cluster-level subsampling used by BrillYurkoWyner2025, who randomly retain a subset of plays per simulated game to reduce within-game dependence when estimating win probability. After sampling, we obtain a total of 2,046 lineout observations for estimation.
@@ 8.2 Limitations and Future Directions
Fourth, selection bias may influence the results (BrillYurkoWyner2025). Stronger teams are both more likely to generate attacking lineouts in the opposition half and more likely to earn penalties. As a result, lineouts in advanced field positions are disproportionately taken by stronger sides, potentially inflating expected points estimates for lineouts near the try line. A similar dynamic is present for the kicking data, since stronger sides will have generated more of the observed penalties, and coaches may only elect to kick when they deem the attempt makeable for their kicker. Future research could focus on further de-biasing EP estimates with respect to the underlying decision processes that generated the data.

### v2

@@ 1 Introduction
While the choice is often straightforward near the opposition’s corner, given a certain point difference, or late in a match, many game states are ambiguous and governed by convention, perceived kicker ability, and the expected quality of a team’s lineout. In contrast, American football has long relied on Expected Points (EP), Expected Points Added (EPA), and win-probability (WP) frameworks to guide fourth-down decision-making, translating field position and context into expected returns and actionable recommendations (Romer, 2006; Burke, 2009; Yurko et al., 2018). In that setting, WP and EP estimators are typically built using statistical machine learning models (e.g., Random Forests and gradient boosting), as is done in Lock and Nettleton (2014) and Baldwin (2021).
In rugby union, public research is only beginning to close this methodological gap, which can be largely attributed to a lack of publicly-available data. Martinez-Arastey et al. (2025) propose a transparent, phase-level modeling framework for calculating Expected Points on English Premiership data, highlighting the difficulty of reliably predicting minority outcomes (e.g., penalty kicks) and providing a reproducible benchmark rather than a deployable decision tool. Separately, public-facing analyses model goal-kicking success as a function of distance and lateral angle, offering practically-useful (albeit context-free) estimates (Monpezat, 2023). Building on these strands—and informed by lessons from the NFL literature on decision modeling—we directly evaluate the penalty decision.
@@ 2.1 Phase-Level Data
Our primary dataset consists of phase-level event logs from the 2018/19 Premiership Rugby season compiled by Martinez-Arastey et al. (2025), covering 132 matches and 35,199 phases of play. Each record describes the on-field state at the start of a phase: the initiating event type, a zonal field location and side, the current score differential, and the next score in points for or against the team in possession. We harmonize timestamps to “seconds remaining in match” so that time is strictly decreasing within a game and directly comparable across halves.
@@ 2.3 Kicking Data
Our kicking data come from the analysis in Quarrie and Hopkins (2015), which studied 582 international rugby matches played between 2002 and 2011. The resulting dataset consists of 3,802 penalties, aggregated into a 5 m-by-5 m grid over the pitch, with the observed proportion of successful kicks recorded for each grid square. These cell-level success rates serve as the raw input for our continuous angle–distance model of kick success. The structure of this dataset is included in Table 2.
@@ 3.2.1 Dealing with Double Counting
This procedure ensures that each run_id contributes at most one observation to the lineout EP model, mitigating bias from repeated, structurally identical possessions. Our approach is conceptually similar to the cluster-level subsampling used by Brill et al. (2025), who randomly retain a subset of plays per simulated game to reduce within-game dependence when estimating win probability. After sampling, we obtain a total of 2,046 lineout observations for estimation.
@@ 8.2 Limitations and Future Directions
Fourth, selection bias may influence the results (Brill et al., 2025). Stronger teams are both more likely to generate attacking lineouts in the opposition half and more likely to earn penalties. As a result, lineouts in advanced field positions are disproportionately taken by stronger sides, potentially inflating expected points estimates for lineouts near the try line. A similar dynamic is present for the kicking data, since stronger sides will have generated more of the observed penalties, and coaches may only elect to kick when they deem the attempt makeable for their kicker. Future research could focus on further de-biasing EP estimates with respect to the underlying decision processes that generated the data.